The average (arithmetic mean) of the measures of angle x, its complementary angle, and its supplementary angle is 80 degrees. what is the measure of the complement of angle x?
It would also help if you could explain how to get the answer.

Answer by lwsshak3(11628)

The average (arithmetic mean) of the measures of angle x, its complementary angle, and its supplementary angle is 80 degrees. what is the measure of the complement of angle x?
It would also help if you could explain how to get the answer.
**
let x=angle
90-x=angle's complement
180-x = angle's supplement
...
average of the three angles as stated:
[x+(90-x)+(180-x)]/3=80
(x+90-x+180-x)/3=80
(-x+270)/3=80
-x+270=3*80=240
-x=-30
x=30
ans:
The measure of the complement of angle x=90-x=90-30=60º